OpenStudy (anonymous):

Find value of 9-24/8+(6*5)

OpenStudy (anonymous):

ok you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I did this but I am getting 24

OpenStudy (anonymous):

Parentheses Exponents Multiplication Division Addition Subtraction

OpenStudy (anonymous):

Please excuse my dear aunt sally

OpenStudy (anonymous):

And the multiple choice has only options 3, 13, 17 and 36

OpenStudy (anonymous):

-15/8 + 30 = 225/8

OpenStudy (anonymous):

9-3+30

hartnn (hartnn):

what is 24/8 = ?

OpenStudy (anonymous):

6+30

OpenStudy (anonymous):

36

OpenStudy (anonymous):

Calle87@ correct

OpenStudy (anonymous):

do parentheses, then division, then left to right for add/subtract

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry my internet stopped working

OpenStudy (anonymous):

for that problem in particular, otherwise use PEMDAS

OpenStudy (anonymous):

Yes I have the PEMDAS written down, thanks.